OpenStudy (anonymous):

2. Jackie, a 25 year old client, want to retire by age 65 with $1,500,000. How much would she have to invest annually, assuming a 6% rate of return?

OpenStudy (anonymous):

2000000 = x * ( r^0 + r^1 + r^2 ... + r^(480-1)) assuming she deposits the money at the end of the month. using the summation of a geometric sequence equation that you learned in junior high and promptly forgot, you get: 2000000 = x * (1 - r^480) / (1 - r) .: x = 2000000 * (1 - 1.004868) / (1 - 1.004868^480) x = $1,048.24 Now if she could get 20% per annum effective on her interest, she would only have to deposit $20.85 per month to meet her goals. She could probably reach her target with as little as $500 a month but given current situations, she should try to put away over $1,000 a month.

OpenStudy (anonymous):

I don't understand where you got your answer from.